a boat moves at a rate of 10mph in still water. if the river's current flows at a rate of 3mph, how long does it take the boat to travel 32 miles upstream?
Answer by eternalblaze21(2)

B is rate of boat=10 mph
W is rate of river/water=3 mph
(B-W)Tu=D <--equation for traveling upstream
(10-3)Tu=32
7Tu=32
Tu=32/7 hrs or 4.6 hrs
Additional notes:
If you were traveling downstream, your equation would look like this: (B+W)Td=D